# Language-Codec: Reducing the Gaps Between Discrete Codec Representation and Speech Language Models

### v3

@@ 3.4. Discriminator and Loss
(10) $\mathcal{L}_{mel}(X,\tilde{X})=\left\|Mel(X)-Mel(\tilde{X})\right\|_{1}$
@@
(11) $\mathcal{L}_{adv}=\frac{1}{K}\sum_{k=1}^{K}max(0,1-D_{k}(\tilde{X}))$
@@
(12) $\mathcal{L}_{feat}=\frac{1}{KL}\sum_{k}\sum_{l}\left\|D_{k}^{l}(X)-D_{k}^{l}(\tilde{X})\right\|_{1}$
@@
(13) $\mathcal{L}_{gen}=\lambda_{q}\mathcal{L}_{q}+\lambda_{mel}\mathcal{L}_{mel}+\lambda_{adv}\mathcal{L}_{adv}+\lambda_{feat}\mathcal{L}_{feat}$
@@
4. Experiments
4.1. Experiment Setup
Datasets
@@ 4.1. Experiment Setup
Language-Codec was trained on a comprehensive 50,000-hour speech dataset. We employed a combination of Librilight’s small and medium collections (Kahn et al., 2020), speech segments from DNS Challenge 4 (Dubey et al., 2022), the Common Voice dataset (version 16.0) (Ardila et al., 2019), LibriTTS (Panayotov et al., 2015) training set, and 20,000 hours of internal Chinese data as the integrated training dataset. To ensure a fair comparison of codec models’ performance, we conducted inference testing on the LibriTTS (Zen et al., 2019) Test-Clean and Test-Other sets to evaluate codecs’ restoration effectiveness in common and noisy environments respectively. Additionally, we performed tests on the LJSpeech dataset to simulate out-of-domain scenarios. For downstream speech language models, we utilized the LibriTTS training set to train zero-shot text-to-speech models. Inference testing was carried out on the LibriSpeech (Panayotov et al., 2015) Test-Clean sets, following VALL-E (Wang et al., 2023) and MobileSpeech (Ji et al., 2024), we filtered audio samples of 4-10 seconds from the LibriSpeech Test-Clean sets.
Baselines
Following Encodec (Défossez et al., 2022), we considered several traditional speech compression models as baselines. Opus (Valin et al., 2012) is a versatile speech and audio codec model that was standardized by the IETF in 2012. EVS (Dietz et al., 2015) is a codec standardized by 3GPP in 2014 and specifically developed for Voice over LTE (VoLTE). We also utilized the official implementation available in Lyra2  [^1] https://github.com/google/lyra at bit rates of 3.2 kbps and 6 kbps. Additionally, we selected three state-of-the-art codec models based on RVQ as baselines. To ensure a fair comparison, we employed the official weight files provided by the Encodec  [^2] https://github.com/facebookresearch/encodec (Défossez et al., 2022) , Vocos  [^3] https://github.com/gemelo-ai/vocos (Siuzdak, 2023) , and SpeechTokenizer  [^4] https://github.com/ZhangXInFD/SpeechTokenizer (Zhang et al., 2023a) frameworks. For downstream speech language models, we replicated two zero-shot TTS models based on discrete codecs modeling: VALL-E (Wang et al., 2023) , representing autoregressive modeling, and MobileSpeech (Ji et al., 2024), representing fully parallel modeling.
Training and Inference Settings
We train Language-Codec up to 2 million iterations, with 1 million iterations allocated to both the generator and discriminator on 8 NVIDIA A100 40G GPUs. Throughout the entire training process, all input speech samples were resampled to 24 kHz, and the batch size was 100. During the training phase, we uniformly truncated excessively long segments in the training data to a fixed length of 10 seconds and subsequently performed a random crop of the waveform to obtain audio snippets of 1-second duration for feeding Language-Codec. Language-Codec is optimized using the AdamW optimizer with an initial learning rate of 2e-4 and betas set to (0.9, 0.999). The learning rate was decayed based on a cosine schedule. MobileSpeech (Ji et al., 2024) was trained for 12 epochs on 8 NVIDIA A100 40G GPUs, with each batch accommodating 3500 frames of the discrete codecs. We optimized the models using the AdamW optimizer with parameters $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95. The learning rate was warmed up for the first 5k updates, reaching a peak of $5\times 10^{-4}$, and then linearly decayed. The AR model and NAR model in VALL-E (Wang et al., 2023) are trained using 4 NVIDIA A100 40GB GPUs with a batch size of 6k acoustic tokens per GPU. We also optimize the models with the AdamW optimizer, warm up the learning rate for the first 5k updates to a peak of $5\times 10^{-4}$, and then linear decay it.
Automatic metrics
@@
For objective evaluation of our codec models, we employ the UTMOS (Saeki et al., 2022) automatic Mean Opinion Score (MOS) prediction system. UTMOS can yield scores highly correlated with human evaluations and is restricted to 16 kHz sample rate. we also adopt the metrics from speech enhancement fields, such as the PESQ (Rix et al., 2001), STOI, and the F1 score for voiced/unvoiced classification (V/UV F1), following the methodology proposed by Vocos(Siuzdak, 2023) to evaluate the performance of discrete codecs. Moreover, we have aligned all our zero-shot TTS experiments metrics with VALL-E (Wang et al., 2023). To evaluate speaker similarity (SPK) between the original prompt and synthesized speech, we employ WavLM-TDNN (Chen et al., 2022). However, due to updates in the repository, we have updated the feature extractor in WavLM, but all our models have been tested by using the same metrics. For assessing automatic speech recognition (ASR) performance, we conduct ASR on the generated audio and calculate the word error rate (WER) compared to the original transcriptions. In this experiment, we utilize the HuBERT-Large (Hsu et al., 2021) model fine-tuned on LibriSpeech 960h as the ASR model. This model is a CTC-based model without language model fusion.
Table 1. The results of different codec models on the LibriTTS Test-Clean and Test-Other dataset.
@@
Model Bandwidth Nq UTMOS $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ V/UV F1 $\uparrow$ SPK $\uparrow$ LibriTTS Test-Clean sets GT - - 4.0562 - - - - Opus 6.0kbps - 2.7961 2.5860 0.9367 0.9408 0.7701 EVS 7.2kbps - 3.4539 3.0988 0.9317 0.9453 0.8524 Lyra-v2 3.2kbps - 2.9856 2.3771 0.8974 0.9438 0.7523 Lyra-v2 6.0kbps - 3.4361 2.9385 0.9257 0.9603 0.8442 Encodec 3.0kbps 4 2.3070 2.0517 0.9007 0.9198 0.7860 Encodec 6.0kbps 8 3.0399 2.7202 0.9391 0.9527 0.8822 Vocos 3.0kbps 4 3.5390 2.4026 0.9231 0.9358 0.7892 Vocos 6.0kbps 8 3.6954 2.8069 0.9426 0.9437 0.8608 SpeechTokenizer 3.0kbps 4 3.5632 1.9311 0.8778 0.9273 0.6587 SpeechTokenizer 6.0kpbs 8 3.8794 2.6121 0.9165 0.9495 0.8311 Language-Codec 3.0kbps 4 3.6191 3.1188 0.9420 0.9618 0.8562 Language-Codec 6.0kbps 8 3.8576 3.6768 0.9609 0.9730 0.9220 LibriTTS Test-Other sets GT - - 3.4831 - - - - Opus 6.0kbps - 2.2628 2.5701 0.9233 0.9265 0.7563 EVS 7.2kbps - 2.8845 2.8456 0.9102 0.9256 0.8407 Lyra-v2 3.2kbps - 2.5187 2.1985 0.8687 0.9146 0.7332 Lyra-v2 6.0kbps - 2.9195 2.7370 0.9012 0.9414 0.8305 Encodec 3.0kbps 4 2.0883 2.0529 0.8835 0.8926 0.7724 Encodec 6.0kbps 8 2.6568 2.6818 0.9241 0.9338 0.8763 Vocos 3.0kbps 4 3.0558 2.1933 0.8967 0.9051 0.7592 Vocos 6.0kbps 8 3.1956 2.5590 0.9209 0.9202 0.8363 SpeechTokenizer 3.0kbps 4 3.0183 1.7373 0.8371 0.8907 0.6071 SpeechTokenizer 6.0kpbs 8 3.2851 2.3269 0.8811 0.9205 0.7925 Language-Codec 3.0kbps 4 3.0202 2.8816 0.9197 0.9465 0.8371 Language-Codec 6.0kbps 8 3.2889 3.5627 0.9461 0.9615 0.9084
Human evaluation
@@
4.2. LibriTTS Evaluation
We evaluated the performance of the codec model on the test set of LibriTTS. The Test-Clean collection consists of 4,837 audio samples, while the Test-Other collection, which mostly contains audio recorded in noisy environments, comprises a total of 5,120 audio samples. Considering that the primary purpose of the discrete codecs is to serve as an audio representation for downstream tasks, excessive channel numbers would significantly burden downstream speech language models. Therefore, we conducted a comparison between four-channel and eight-channel dimensions. Among the objective metrics we employed, UTMOS and speaker similarity metrics closely approximate the subjective perception of human listeners. On the other hand, PESQ, STOI, and F1 metrics are more indicative of the inherent quality of the audio signal.
Based on the observations from Table 1, the following conclusions can be drawn: 1) Regarding the audio reconstruction of the four-channel codecs, the Language-Codec model significantly outperforms all baseline models in terms of objective metrics. While there is a slight decrease in audio reconstruction quality when the number of channels is reduced from eight to four in the baseline models, the Language-Codec model maintains a consistently good reconstruction performance. Additionally, it is noteworthy that the four-channel reconstruction of Language-Codec even surpasses the eight-channel performance of several baseline models. For instance, in terms of the PESQ and STOI metrics, the four-channel Language-Codec model outperforms the eight-channel SpeechTokenizer model by 0.03 and 0.5 in LibriTTS Test-Clean sets. Furthermore, in the UTMOS metric, the four-channel Language-Codec model significantly outperforms the eight-channel Encodec model. 2) In the eight-channel codecs reconstruction, the Language-Codec model also maintains SOTA reconstruction quality. Although the eight-channel SpeechTokenizer model achieved similar scores to the Language-Codec model in terms of the UTMOS metric, it significantly underperformed in other metrics such as SOTI, SPK, and PESQ compared to the Language-Codec model, and even performed noticeably worse than the Encodec model. Considering the overall auditory perception and average audio quality, the Language-Codec model achieves the best performance. 3) We noticed that all comparative models maintain similar conclusions and trends between the Test-Clean (clean dataset) and Test-Other (noisy dataset) conditions. Moreover, the Language-Codec model demonstrates good reconstruction quality even in noisy environments. 4) It is worth mentioning that the commonly employed Encodec model in downstream tasks consistently performs lower than the Language-Codec model in the UTMOS metric. Upon carefully listening to relevant audio samples, we identified that the Encodec model may introduce more reconstruction artifacts, a characteristic that significantly affects UTMOS scores.
Table 2. The results of different codec models on the LJSpeech dataset.
Model Bandwidth Nq UTMOS $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ V/UV F1 $\uparrow$ SPK $\uparrow$ LJSpeech GT - - 4.3794 - - - - Opus 6.0kbps - 2.7640 2.1433 0.1245 0.9489 0.7098 EVS 7.2kbps - 3.8991 3.0560 0.9507 0.9521 0.8551 Lyra-v2 3.2kbps - 3.3773 2.4182 0.9161 0.9421 0.7041 Lyra-v2 6.0kbps - 3.9591 2.8853 0.9418 0.9551 0.8007 Encodec 3.0kbps 4 2.3905 2.0194 0.9058 0.9326 0.8177 Encodec 6.0kbps 8 3.2286 2.6633 0.9441 0.9555 0.8952 Vocos 3.0kbps 4 3.7880 2.5006 0.9310 0.9388 0.7801 Vocos 6.0kbps 8 4.0332 2.9258 0.9497 0.9459 0.8339 SpeechTokenizer 3.0kbps 4 3.9908 2.0458 0.9021 0.9299 0.6793 SpeechTokenizer 6.0kpbs 8 4.2373 2.6413 0.9316 0.9452 0.8332 Language-Codec 3.0kbps 4 3.9461 3.0772 0.9468 0.9573 0.8591 Language-Codec 6.0kbps 8 4.2221 3.6186 0.9662 0.9676 0.9269
4.3. Generalization Evaluation

### v4

@@ 3.4 Discriminator and Loss
$\mathcal{L}_{mel}(X,\tilde{X})=\left\|Mel(X)-Mel(\tilde{X})\right\|_{1}$ (10)
@@
$\mathcal{L}_{adv}=\frac{1}{K}\sum_{k=1}^{K}max(0,1-D_{k}(\tilde{X}))$ (11)
@@
$\mathcal{L}_{feat}=\frac{1}{K*L}\sum_{k}\sum_{l}\left\|D_{k}^{l}(X)-D_{k}^{l}(\tilde{X})\right\|_{1}$ (12)
@@
$\mathcal{L}_{gen}=\lambda_{q}\mathcal{L}_{q}+\lambda_{mel}\mathcal{L}_{mel}+\lambda_{adv}\mathcal{L}_{adv}+\lambda_{feat}\mathcal{L}_{feat}$ (13)
@@
4 Experiments
4.1 Experiment Setup
@@
Datasets. Language-Codec was trained on a comprehensive 50,000-hour speech dataset (We verify the effect of training dataset of different sizes on the model in Appendix A). We employed a combination of Librilight’s small and medium collections Kahn et al. (2020), speech segments from DNS Challenge 4 Dubey et al. (2022), the Common Voice dataset (version 16.0) Ardila et al. (2019), LibriTTS Panayotov et al. (2015) training set, and 20,000 hours of internal Chinese data as the integrated training dataset. To ensure a fair comparison of codec models’ performance, we conducted inference testing on the LibriTTS Zen et al. (2019) Test-Clean and Test-Other sets to evaluate codecs’ restoration effectiveness in common and noisy environments respectively. Additionally, we performed tests on the LJSpeech dataset to simulate out-of-domain scenarios. For downstream speech language models, we utilized the LibriTTS training set to train zero-shot text-to-speech models. Inference testing was carried out on the LibriSpeech Panayotov et al. (2015) Test-Clean sets, following VALL-E Wang et al. (2023) and MobileSpeech Ji et al. (2024a), we filtered audio samples of 4-10 seconds from the LibriSpeech Test-Clean sets.
@@
Automatic metrics. For objective evaluation of our codec models, we employ the UTMOS Saeki et al. (2022) automatic Mean Opinion Score (MOS) prediction system. UTMOS can yield scores highly correlated with human evaluations and is restricted to 16 kHz sample rate. we also adopt the metrics from speech enhancement fields, such as the PESQ Rix et al. (2001), STOI, and the F1 score for voiced/unvoiced classification (V/UV F1), following the methodology proposed by VocosSiuzdak (2023) to evaluate the performance of discrete codecs. Moreover, we have aligned all our zero-shot TTS experiments metrics with VALL-E Wang et al. (2023). To evaluate speaker similarity (SPK) between the original prompt and synthesized speech, we employ WavLM-TDNN Chen et al. (2022). However, due to updates in the repository, we have updated the feature extractor in WavLM, but all our models have been tested by using the same metrics. For assessing automatic speech recognition (ASR) performance, we conduct ASR on the generated audio and calculate the word error rate (WER) compared to the original transcriptions. In this experiment, we utilize the HuBERT-Large Hsu et al. (2021) model fine-tuned on LibriSpeech 960h as the ASR model. This model is a CTC-based model without language model fusion.
Due to space constraints, details of the Baselines and Training and Inference Settings are in Appendix B and Appendix C. Subjective evaluations are detailed in Appendix D.
Table 1: The results of different codec models on the LibriTTS Test-Clean and Test-Other dataset. Since downstream generative models use at most 8 layers of quantizers, we limit our comparisons to the 6 kbps and 3 kbps dimensions, even though 12 kbps can achieve better reconstruction performance.
@@
Model Bandwidth Nq UTMOS $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ V/UV F1 $\uparrow$ SPK $\uparrow$ LibriTTS Test-Clean sets GT - - 4.0562 - - - - Opus 6.0kbps - 2.7961 2.5860 0.9367 0.9408 0.7701 EVS 7.2kbps - 3.4539 3.0988 0.9317 0.9453 0.8524 Encodec 3.0kbps 4 2.3070 2.0517 0.9007 0.9198 0.7860 Encodec 6.0kbps 8 3.0399 2.7202 0.9391 0.9527 0.8822 Vocos 3.0kbps 4 3.5390 2.4026 0.9231 0.9358 0.7892 Vocos 6.0kbps 8 3.6954 2.8069 0.9426 0.9437 0.8608 SpeechTokenizer 3.0kbps 4 3.5632 1.9311 0.8778 0.9273 0.6587 SpeechTokenizer 6.0kpbs 8 3.8794 2.6121 0.9165 0.9495 0.8311 DAC 3.0kbps 4 2.9902 2.4091 0.9118 0.9531 0.8129 DAC 6.0kbps 8 3.6804 3.5558 0.9257 0.9711 0.8715 Language-Codec 3.0kbps 4 3.7875 3.2675 0.9493 0.9657 0.8698 Language-Codec 6.0kbps 8 4.0372 3.8813 0.9715 0.9754 0.9318
4.2 Reconstruction Evaluation
We evaluated the reconstruction performance of the codec model on the LibriTTS Test-Clean dataset Zen et al. (2019). Considering that the primary purpose of the discrete codecs is to serve as an audio representation for downstream tasks, excessive channel numbers would significantly burden downstream speech language models. Therefore, we conducted a comparison between four-channel and eight-channel dimensions. We believe that the results on the LibriTTS Test-Clean subset adequately reflect the reconstruction performance of the codec model. To further evaluate the reconstruction performance of the codec model in noisy environments and out-of-domain scenarios, we additionally compare the performance of different codec models on the LibriTTS Test-Other Zen et al. (2019) and LJSpeech datasets in Appendix E.
Based on the observations from Table 1, the following conclusions can be drawn: 1) Regarding the audio reconstruction of the four-channel codecs, the Language-Codec model significantly outperforms all baseline models in terms of objective metrics. While there is a slight decrease in audio reconstruction quality when the number of channels is reduced from eight to four in the baseline models, the Language-Codec model maintains a consistently good reconstruction performance. Additionally, it is noteworthy that the four-channel reconstruction of Language-Codec even surpasses the eight-channel performance of several baseline models. For instance, in terms of the PESQ and STOI metrics, the four-channel Language-Codec model outperforms the eight-channel SpeechTokenizer model by 0.6 and 0.03 in LibriTTS Test-Clean sets. Furthermore, in the UTMOS metric, the four-channel Language-Codec model significantly outperforms the eight-channel Encodec model. 2) In the eight-channel codecs reconstruction, the Language-Codec model also maintains SOTA reconstruction quality. Although the eight-channel SpeechTokenizer model achieved similar scores to the Language-Codec model in terms of the UTMOS metric, it significantly underperformed in other metrics such as SOTI, SPK, and PESQ compared to the Language-Codec model, and even performed noticeably worse than the Encodec model. Considering the overall auditory perception and average audio quality, the Language-Codec model achieves the best performance.
4.3 Zero-shot TTS Evaluation
Regarding the inference phase, to ensure fair comparisons, we followed the experimental protocols outlined in VALL-E Wang et al. (2023) and employed the LibriSpeech Test-Clean dataset Panayotov et al. (2015), ensuring no overlap with our training data. We specifically utilized samples from the LibriSpeech Test-Clean dataset with durations ranging from 4 to 10 seconds, resulting in a subset of 2.2 hours. Following VALL-E, we use the whole transcription and the first 3 seconds of the utterance as the phoneme and acoustic prompts respectively, and ask the model to generate the continuations.
Table 2: Evaluation of zero-shot TTS models with different codecs on the LibriSpeech Test-Clean corpus.
Zero-shot TTS Model WER $\downarrow$ SPK $\uparrow$ MOS-Q $\uparrow$ MOS-P $\uparrow$ MOS-S $\uparrow$ VALL-E w/ Encodec 4.3 0.6115 3.73$\pm$0.09 3.76$\pm$0.12 3.74$\pm$0.11 VALL-E w/ Encodec+Vocos 4.3 0.6198 3.83$\pm$0.07 3.81$\pm$0.13 3.82$\pm$0.06 VALL-E w/ Language-Codec 3.8 0.6995 4.01$\pm$0.09 3.89$\pm$0.12 3.99$\pm$0.08 MobileSpeech w/ Encodec 3.2 0.6776 3.91$\pm$0.10 3.99$\pm$0.12 3.98$\pm$0.11 Mobilespeech w/ Encodec+Vocos 3.1 0.6883 4.05$\pm$0.08 4.03$\pm$0.12 4.04$\pm$0.09 MobileSpeech w/ Language-Codec 2.9 0.7712 4.20$\pm$0.11 4.09$\pm$0.07 4.18$\pm$0.10
Given the widespread adoption of Encodec models in downstream speech language models for extracting features as intermediate acoustic tokens, we have selected it as the primary baseline model in our study. For the Vocos version, we maintain the practice of training downstream models using feature representations extracted by the Encodec model. However, during the inference and decoding phase, we replace the Encodec decoder with the Vocos decoder. A more comprehensive explanation of this approach, including the rationale and methodology behind it, will be provided in the subsequent section dedicated to ablation experiments. Additionally, in the case of the Language-Codec version, we focus exclusively on training models with codec representations derived from the Language-Codec model.
[…]
4.4 Ablation experiment
In this section, we conducted a detailed analysis of MCRVQ module. The ablation experiments regarding the new decoder and the multi-scale discriminator are presented in Appendix G and F.
We validated the role of the Masked Channel Residual Vector Quantization (MCRVQ) module in the language-codec model. Considering that the design purpose of the MCRVQ mechanism is to reduce the difficulty of text generation in downstream tasks, we conducted ablation experiments on the zero-shot TTS model downstream. Specifically, we first replaced the MCRVQ module with the RVQ module (eight quantizers) while keeping the same training steps and other configurations. We refer to this experiment setup as Language-Codec w/o MCRVQ. We used Language-Codec w/o MCRVQ to extract the corresponding discrete codec features and retrained the downstream VALL-E and MobileSpeech models. The experimental results, as shown in Table 3, revealed that there was no significant difference (0.2) between Language-Codec w/o MCRVQ and Language-Codec in terms of the robustness metric WER. However, in terms of the objective metric of speaker similarity, omitting the MCRVQ module resulted in a decrease of 0.06 similarity in VALL-E and MobileSpeech, respectively, indicating that the MCRVQ module indeed enhances the codec generation capability of the downstream speech synthesis model by weakening the difficulty of text generation for codec.
In addition, we also conducted corresponding subjective CMOS tests. From Table 4, it can be observed that in the autoregressive discrete codec modeling experiments of the VALL-E model, the CMOS values of the synthesized audio decreased by 0.19 when the MCRVQ module was omitted compared to the original Language-Codec model. Similarly, in the parallel discrete codec modeling experiments of the MobileSpeech model, the CMOS values of the synthesized audio decreased by 0.25 when the MCRVQ module was omitted compared to the original Language-Codec model, which further indicated that the codec generated by the Language-Codec w/o MCRVQ model had lower subjective audio quality and audio similarity than the codec generated by the Language-Codec.
Table 3: The ablation experiments of the MCRVQ module, we assessed the performance of WER and SPK metrics.
Model Codec Model WER $\downarrow$ SPK $\uparrow$ VALL-E Language-Codec w/o MCRVQ 4.1 0.6383 VALL-E Language-Codec 3.8 0.6995 MobileSpeech Language-Codec w/o MCRVQ 3.1 0.7103 MobileSpeech Language-Codec 2.9 0.7712
Table 4: The ablation experiments of the MCRVQ module, we assessed the performance of subjective metric CMOS.
Model Codec Model CMOS $\uparrow$ VALL-E Language-Codec w/o MCRVQ -0.19 VALL-E Language-Codec 0.00 MobileSpeech Language-Codec w/o MCRVQ -0.25 MobileSpeech Language-Codec 0.00
5 Conclusions
In this article, we propose Language-Codec, a discrete acoustic codec model that enhances adaptation to downstream speech language models. Through improved model architecture and a unique Masked Channel residual vector quantization mechanism, we achieve excellent audio reconstruction quality with just four layers of codecs. The Language-Codec model demonstrates effective audio restoration performance in both clean audio and noisy environments. Furthermore, we validate the generalization capability of Language-Codec in unseen domains and its generation ability in downstream zero-shot TTS models, yielding satisfactory results. We envision Language-Codec as a state-of-the-art foundational codec model for future research in the field of speech generation.
6 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No.62222211 and No.U24A20326
7 Limitations
While Language-Codec models demonstrate superior reconstruction quality compared to some notable comparative models, it is important to note that due to time constraints, Language-Codec models have been trained solely on speech corpora and validated solely on downstream speech language models. Although the majority of codec models also exclusively support speech data, we aspire to develop a more comprehensive and versatile codec model. In the future, we plan to incorporate a larger training dataset (consisting of several hundred thousand hours) to encompass a wider range of signal types, including audio, music, and more.
Appendix A Ablation experiments about varied training data volumes
As shown in Table 5. We found that the size of the dataset does not significantly impact speech reconstruction. Particularly noteworthy is the internal dataset comprising 20k hours, which, given its Chinese language corpus, exerts minimal influence on the results presented in the paper. Dataset size may more influence generalization (more languages) instead of reconstruction.
Appendix B Baselines
Following Encodec Défossez et al. (2022), we considered several traditional speech compression models as baselines. Opus Valin et al. (2012) is a versatile speech and audio codec model that was standardized by the IETF in 2012. EVS Dietz et al. (2015) is a codec standardized by 3GPP in 2014 and specifically developed for Voice over LTE (VoLTE). We also utilized the official implementation available in Lyra2  [^1] https://github.com/google/lyra at bit rates of 3.2 kbps and 6 kbps. Additionally, we selected three state-of-the-art codec models based on RVQ as baselines. To ensure a fair comparison, we employed the official weight files provided by the Encodec  [^2] https://github.com/facebookresearch/encodec Défossez et al. (2022) , Vocos  [^3] https://github.com/gemelo-ai/vocos Siuzdak (2023) , SpeechTokenizer  [^4] https://github.com/ZhangXInFD/SpeechTokenizer Zhang et al. (2023) and DAC  [^5] https://github.com/descriptinc/descript-audio-codec Kumar et al. (2024a) frameworks. For downstream speech language models, we replicated two zero-shot TTS models based on discrete codecs modeling: VALL-E Wang et al. (2023) , representing autoregressive modeling, and MobileSpeech Ji et al. (2024a), representing fully parallel modeling.
Appendix C Training and Inference Settings
We train Language-Codec up to 2 million iterations, with 1 million iterations allocated to both the generator and discriminator on 8 NVIDIA A100 40G GPUs. Throughout the entire training process, all input speech samples were resampled to 24 kHz, and the batch size was 100. During the training phase, we uniformly truncated excessively long segments in the training data to a fixed length of 10 seconds and subsequently performed a random crop of the waveform to obtain audio snippets of 1-second duration for feeding Language-Codec. Language-Codec is optimized using the AdamW optimizer with an initial learning rate of 2e-4 and betas set to (0.9, 0.999). The learning rate was decayed based on a cosine schedule. MobileSpeech Ji et al. (2024a) was trained for 12 epochs on 8 NVIDIA A100 40G GPUs, with each batch accommodating 3500 frames of the discrete codecs. We optimized the models using the AdamW optimizer with parameters $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95. The learning rate was warmed up for the first 5k updates, reaching a peak of $5\times 10^{-4}$, and then linearly decayed. The AR model and NAR model in VALL-E Wang et al. (2023) are trained using 4 NVIDIA A100 40GB GPUs with a batch size of 6k acoustic tokens per GPU. We also optimize the models with the AdamW optimizer, warm up the learning rate for the first 5k updates to a peak of $5\times 10^{-4}$, and then linear decay it.
Appendix D Human evaluation
@@
Appendix E More Reconstruction Evaluation
We evaluate the performance of different codec models in noisy environments using the LibriTTS Test-Other dataset. The experimental results are presented in Table 6. We noticed that all comparative models maintain similar conclusions and trends between the Test-Clean (clean dataset) and Test-Other (noisy dataset) conditions. Moreover, the Language-Codec model demonstrates good reconstruction quality even in noisy environments.